OpenStudy (solomonzelman):

@laurenlauren12345, here!

OpenStudy (anonymous):

Hi !!! What is the slope intercept form that passes through the given point and is perpendicular (-6,4), 3y =2x-3? What is the answer?! Help!! :)

OpenStudy (solomonzelman):

Slopes y-intercepts: \[\huge\color{blue}{ y=mx+b } \] Slope intercept form.

OpenStudy (solomonzelman):

Lets take an example:\[\huge\color{blue}{ y=3x+2 } \] how would you know how to graph the line? choose any x-value. lets make x equal -1, plug it in, \[\huge\color{blue}{ y=3x+2 } \]\[\huge\color{blue}{ y=3(-1)+2 } \]\[\huge\color{blue}{ y=-3+2 } \]\[\huge\color{blue}{ y=-1 } \]

OpenStudy (solomonzelman):

So we know that when x=-1 y=-1 (y and x don't have to be the same, this just happened like this), so your point (one of them, would be (-1,-1)

OpenStudy (anonymous):

Wut is the answer?

OpenStudy (solomonzelman):

perpendicular to 3y =2x-3 line, and passes through (-6,4) write your equation in y=mx+b form, divide everything by 3, 3y/3=2x/2-2 y=2/3 x - 2/3 perpendicular slope to any slope (let slope represent m, is) -1/m.

OpenStudy (solomonzelman):

-3/2 would be perpendicular to 2/3

OpenStudy (anonymous):

So that is the answer

OpenStudy (solomonzelman):

So, now, the y-intercept is unknown, say, y= -2/3 x +b plug in your point, 4=-2/3 (6) + b 4 = -4 +b b=8 KNOWING your y-intercept=8 and slope =-3/2 the answer is?

OpenStudy (anonymous):

Y=-3/2 +8

OpenStudy (solomonzelman):

Close, y=-3/2x+8

OpenStudy (solomonzelman):

as y=mx+b